OpenStudy (anonymous):

if f(x)=1/(x+2), what is f(3+h)? Is the answer 1/(h+5)?

OpenStudy (anonymous):

\[f(x) = 1/x+2\] f(3+h) = 1/ [(3+h) + 2] f(3+h) = 1/(h+5) Correct.

OpenStudy (anonymous):

thanks again :P

OpenStudy (anonymous):

=0)